# Logarithmic Smoothing for Adaptive PAC-Bayesian Off-Policy Learning

### v1

@@ F.1 Detailed Setup
Our experimental framework follows the standard multiclass-to-bandit conversion widely used in prior work [15, 47]. Each multiclass dataset consists of feature-label pairs, which we transform into contextual bandit problems by treating features as contexts and labels as actions. The reward $r$ for taking action $a$ for context $x$ is modeled as Bernoulli with probability $p_{x}=\epsilon+\mathds{1}\left[a=a^{\star}(x)\right](1-2\epsilon)$, where $a^{\star}(x)$ is the true label of features $x$, and $\epsilon$ is a noise parameter, that we set to $\epsilon=0.2$ in all our experiments. This setup ensures an average reward of $1-\epsilon$ for the optimal action $a^{\star}(x)$ and $\epsilon$ for all others. In our experiments, we use parametrized Linear Gaussian Policies [41, 3], these policies interact with the contextual bandit environment to construct a logged bandit feedback dataset in the form $\{x_{i},a_{i},c_{i}\}_{i\in[n]}$, where $c_{i}=-r_{i}$ is the associated cost.
@@
All datasets are used without processing aside from CIFAR100 that we embed using a pretrained Resnet, obtaining a low dimensional vector of $100$ for each image.
Table 2: Detailed statistics of the datasets used.
Datasets $N$ $n_{l}$ $n_{c}$ $n_{test}$ $K$ $p$
@@
We use the Linear Gaussian Policy of [41]. To obtain these policies, we restrict $f_{\theta}$ to:
@@
$\pi_{\boldsymbol{\mu},\sigma}(a|x)$ $=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,1)}\left[\prod_{a^{\prime}\neq a}\Phi\left(\epsilon+\frac{\phi(x)^{T}(\boldsymbol{\mu}_{a}-\boldsymbol{\mu}_{a^{\prime}})}{\sigma||\phi(x)||}\right)\right]$
@@
$\pi_{0}$ is trained on $D_{l}$ (supervised manner) with the following parameters: We use $L_{2}$ regularization of $10^{-4}$. This is used to prevent the logging policy $\pi_{0}$ from being close to deterministic, allowing efficient learning with importance sampling. We use Adam [29] with a learning rate of $10^{-1}$ for $10$ epochs. Once it is trained, we use an inverse temperature parameter $\alpha$ on its score to interpolate between a uniform policy $\alpha=0$ and a trained policy $\alpha=1$.
@@
In each optimization subroutine, we use Adam [29] with a learning rate of $10^{-3}$ for $10$ epochs. The gradient of LIG policies is a one dimensional integral, and is approximated using $S=32$ samples.
$\pi_{\boldsymbol{\mu},\sigma}(a|x)$ $=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,1)}\left[\prod_{a^{\prime}\neq a}\Phi\left(\epsilon+\frac{\phi(x)^{T}(\boldsymbol{\mu}_{a}-\boldsymbol{\mu}_{a^{\prime}})}{\sigma||\phi(x)||}\right)\right]$
$\approx\frac{1}{S}\sum_{s=1}^{S}\prod_{a^{\prime}\neq a}\Phi\left(\epsilon_{s}+\frac{\phi(x)^{T}(\boldsymbol{\mu}_{a}-\boldsymbol{\mu}_{a^{\prime}})}{\sigma||\phi(x)||}\right)\quad\epsilon_{1},...,\epsilon_{S}\sim\mathcal{N}(0,1).$
F.2 More Experiments
@@
In the main experiments, we used $\alpha=0.2$, which makes the starting behavior policy keep some information about the good actions while being diffuse enough to put mass on all actions. We give here a more comprehensive view of how the algorithms behave depending on $\alpha$. The main take away, which reflects our first intuition, is that if we start far from $\pi_{\star}$, the improvement brought by adaptive off-policy approaches is substantial compared to using the static behavior policy $\pi_{0}$ to collect data.
Effect of Multiple Deployments while varying $\alpha$. We focus on CIFAR100 of large action space ($|\mathcal{A}|=100$). $\alpha=0.$ is the purely uniform behavior policy $\pi_{0}$, $\alpha=0.5$ is a very good policy but is still diffuse and $\alpha=1.$ is a peaked policy. Results are shown in Table 3. adaAdjLS outperforms adaLS in all scenarios, with a bigger gap when $\pi_{0}$ is far from a good policy.
@@
$\alpha$ $k$ adaLS (Alg 1) adaAdjLS (Alg 2)
$\alpha=0.$ 00 $-0.010$ $-0.010$
$1$ $-0.011$ $-0.011$
$5$ $-0.011$ $-0.011$
$10$ $-0.020$ $-0.026$
$100$ $-0.044$ $\mathbf{-0.077}$
$\alpha=0.5$ 00 $-0.268$ $-0.268$
$1$ $-0.430$ $-0.440$
$5$ $-0.500$ $-0.520$
$10$ $-0.527$ $-0.555$
$100$ $-0.583$ $\mathbf{-0.621}$
$\alpha=1.$ 00 $-0.531$ $-0.531$
$1$ $-0.569$ $-0.572$
$5$ $-0.590$ $-0.600$
$10$ $-0.600$ $-0.614$
$100$ $-0.630$ $\mathbf{-0.649}$
Effect of $\alpha$ on SCRM. We extend our evaluation of SCRM by varying the parameter $\alpha\in{0,0.5,1}$, with $k$ fixed at 10 across all datasets. When $\alpha=0$, the behavior policy is entirely uniform, making the learning task considerably more difficult—particularly in environments with large action spaces—since the policy provides minimal guidance and infrequently selects low-cost actions. Despite these challenges, both adaLS and adaAdjLS consistently outperform SCRM. As $\alpha$ increases, the behavior policy becomes more informative, improving SCRM’s performance. Nonetheless, our adaptive methods, especially adaAdjLS, continue to achieve superior results. The accompanying plots highlight the consistent advantage of adaAdjLS across all settings.

### v2

@@ E.1 Detailed Setup
Our experimental framework follows the standard multiclass-to-bandit conversion widely used in prior work [Dudik et al., 2014, Swaminathan and Joachims, 2015]. Each multiclass dataset consists of feature-label pairs, which we transform into contextual bandit problems by treating features as contexts and labels as actions. The reward $r$ for taking action $a$ for context $x$ is modeled as Bernoulli with probability $p_{x}=\epsilon+\mathds{1}\left[a=a^{\star}(x)\right](1-2\epsilon)$, where $a^{\star}(x)$ is the true label of features $x$, and $\epsilon$ is a noise parameter, that we set to $\epsilon=0.2$ in all our experiments. This setup ensures an average reward of $1-\epsilon$ for the optimal action $a^{\star}(x)$ and $\epsilon$ for all others. In our experiments, we use parametrized Linear Gaussian Policies [Sakhi et al., 2023a, Aouali et al., 2023], these policies interact with the contextual bandit environment to construct a logged bandit feedback dataset in the form $\{x_{i},a_{i},c_{i}\}_{i\in[n]}$, where $c_{i}=-r_{i}$ is the associated cost.
@@
All datasets are used without processing aside from CIFAR100 that we embed using a pretrained Resnet, obtaining a low dimensional vector of $100$ for each image. All experiments do not require heavy computations, they were conducted on a single 16 CPU/ 1 GPU V100 machine.
Datasets NN $n_{l}$ $n_{c}$ $n_{test}$ $K$ $p$
@@
We use the Linear Gaussian Policy of [Sakhi et al., 2023a]. To obtain these policies, we restrict $f_{\theta}$ to:
@@
[…]
@@
$\pi_{0}$ is trained on $D_{l}$ (supervised manner) with the following parameters: We use $L_{2}$ regularization of $10^{-4}$. This is used to prevent the logging policy $\pi_{0}$ from being close to deterministic, allowing efficient learning with importance sampling. We use Adam [Kingma and Ba, 2014] with a learning rate of $10^{-1}$ for $10$ epochs. Once it is trained, we use an inverse temperature parameter $\alpha$ on its score to interpolate between a uniform policy $\alpha=0$ and a trained policy $\alpha=1$.
@@
In each optimization subroutine, we use Adam with a learning rate of $10^{-3}$ for $10$ epochs. The gradient of LIG policies is a one dimensional integral, and is approximated using $S=32$ samples.
[…]
E.2 More Experiments
@@
In the main experiments, we used $\alpha=0.2$, which makes the starting behavior policy keep some information about the good actions while being diffuse enough to put mass on all actions. We give here a more comprehensive view of how the algorithms behave depending on $\alpha$. The main take away, which reflects our first intuition, is that if we start far from $\pi_{\star}$, the improvement brought by sequential off-policy approaches is substantial compared to using the static behavior policy $\pi_{0}$ to collect data.
Effect of Multiple Deployments while varying $\alpha$. We focus on CIFAR100 of large action space ($|\mathcal{A}|=100$). $\alpha=0.$ is the purely uniform behavior policy $\pi_{0}$, $\alpha=0.5$ is a very good policy but is still diffuse and $\alpha=1.$ is a peaked policy. Results are shown in Table˜3. SeqAdjLS outperforms SeqLS in all scenarios, with a bigger gap when $\pi_{0}$ is far from a good policy.
$\alpha$ $k$ SeqLS (Alg 1) SeqAdjLS (Alg 2)
$\alpha=0.$ 0 0.010-0.010 0.010-0.010
11 0.011-0.011 0.011-0.011
55 0.011-0.011 0.011-0.011
1010 0.020-0.020 0.026-0.026
100100 0.044-0.044 $\mathbf{-0.077}$
$\alpha=0.5$ 0 0.268-0.268 0.268-0.268
11 0.430-0.430 0.440-0.440
55 0.500-0.500 0.520-0.520
1010 0.527-0.527 0.555-0.555
100100 0.583-0.583 $\mathbf{-0.621}$
$\alpha=1.$ 0 0.531-0.531 0.531-0.531
11 0.569-0.569 0.572-0.572
55 0.590-0.590 0.600-0.600
1010 0.600-0.600 0.614-0.614
100100 0.630-0.630 $\mathbf{-0.649}$
@@
Effect of $\alpha$ on SCRM. We extend our evaluation of SCRM by varying the parameter $\alpha\in{0,0.5,1}$, with $k$ fixed at 10 across all datasets. When $\alpha=0$, the behavior policy is entirely uniform, making the learning task considerably more difficult—particularly in environments with large action spaces—since the policy provides minimal guidance and infrequently selects low-cost actions. Despite these challenges, both SeqLS and SeqAdjLS consistently outperform SCRM. As $\alpha$ increases, the behavior policy becomes more informative, improving SCRM’s performance. Nonetheless, our sequential methods, especially SeqAdjLS, continue to achieve superior results. The accompanying plots highlight the consistent advantage of SeqAdjLS across all settings.